The Malayo Colleges Mindanao recently conducted a study analyzing the privacy management habits of teen internet users. In a group of 173 teens, 139 reported having more than 100 followers on Twitter. Find the margin of error of a 90% confidence interval for the true proportion of teens who would report having more than 100 followers on Twitter.



1
Expert's answer
2022-03-08T12:44:13-0500

"p=\frac{139}{173}=0.8035."

"ME=1.645\sqrt{\frac{0.8035(1-0.8035)}{173}}=0.0497."